OpenStudy (anonymous):

Find the equation of a straight line that passes through: (2, 1) and (-2, -7) Using: y2-y1 m=------ x2-x1 Can anyone help me by verifying if is this the correct answer? Thanks

OpenStudy (anonymous):

y = -8x + b 1 = -8(1) + b 1 = -8 +b 8 + 1 = b 9 = b y = -8x + 9

OpenStudy (anonymous):

\[y = mx +c\] \[m = 2\] \[c =-3\] \[y =2x - 3\]